OpenStudy (anonymous):

The area of a rectangle 12 feet by 18 feet is?

OpenStudy (saifoo.khan):

16 x 18 = 288 ft.sq.

myininaya (myininaya):

12*18=216 square ft

OpenStudy (saifoo.khan):

216* sorry!

myininaya (myininaya):

lol

OpenStudy (anonymous):

my answer choices are A)6 sq yards B)8 sq yrds C)24 sq yrds D)36 sq yrds

myininaya (myininaya):

convert it to sq yards then

myininaya (myininaya):

1 yard =3 feet (1 yard)^2=(3 feet)^2 1 square yard=9 square feet if we got 216 square feet, then 216/9*(1square yard)=216/9*(9 square feet) 216/9 square yard=216 square feet

myininaya (myininaya):

24 square yard=216 square feet

OpenStudy (anonymous):

ok thank you so much

OpenStudy (saifoo.khan):

crrct.